OpenStudy (anonymous):

good morning everyone!!!!! can someone please help me with Algebraic Expressions I think i know how but there a variables throwing me off course... Thank you

OpenStudy (anonymous):

this is the first exspression im caught on http://static.k12.com/bank_packages/files/media/mathml_3ad92e4b86aec91607782447dc174d5c7d8ad588_1.gif

OpenStudy (anonymous):

what should i hv to do

OpenStudy (anonymous):

i just need someone to explain how i would use the variable x in the answer... i don't understand the problem

OpenStudy (anonymous):

@jojo4eva

OpenStudy (anonymous):

@Vivid

OpenStudy (anonymous):

@Legends

OpenStudy (a_clan):

We cannot subtract a constant from a variable. So, you will have to leave the answer in terms of 'x'; but remaining expression has to be simplified to lowest possible form.

OpenStudy (anonymous):

ok i unsterstand the first part... but what do you mean by the second?

OpenStudy (anonymous):

@A_clan

OpenStudy (anonymous):

@k12andstudyislandhelp

OpenStudy (a_clan):

Have you heard of BODMAS rule?

OpenStudy (anonymous):

no i havent... what is it? @A_clan

OpenStudy (a_clan):

B → Brackets first (parentheses) O → Of (orders i.e. Powers and Square Roots, Cube Roots, etc.) DM → Division and Multiplication (start from left to right) AS → Addition and Subtraction (start from left to right) Note: (i) Start Divide/Multiply from left side to right side since they perform equally. (ii) Start Add/Subtract from left side to right side since they perform equally. ( I searched it on GOOGLE.)

OpenStudy (anonymous):

oh ok thank you... so i figured out the the parentheses which was easy but now i dont know wht to do with the 2x!

OpenStudy (anonymous):

@A_clan

OpenStudy (a_clan):

This is the priority rule. Whenever you encounter a mathematical expression containing one or more of mathematical operations BODMAS, then start solving the brackets first , followed by 'Off' , then DM and finally AS. Within the 'Brackets' also start by solving bar bracket first, followed by small brackets, then parentheses and finally square brackets.

OpenStudy (a_clan):

In the given question, we shall start by solving the small brackets. (0.5- 1.4) =(-0.9)

OpenStudy (anonymous):

then you subtract that with 2x

OpenStudy (anonymous):

but i thought i couldnt do that because of the variable?

OpenStudy (a_clan):

Next step is the square bracket. [2x- (-0.9)] = [2x+0.9]

OpenStudy (anonymous):

ahh i would have never thought to do that lol

OpenStudy (anonymous):

then after you do this you multiply that by 4.5

OpenStudy (anonymous):

13.05?

OpenStudy (a_clan):

Then in last step multiply this square bracket. -4.5[2x+0.9] =-4.5 * 2x + -4.5*0.9 =-9x - 4.05

OpenStudy (a_clan):

This is the most simplified form. and hence the required answer.

OpenStudy (anonymous):

thank you!!!! i think i get it now!!!

OpenStudy (a_clan):

you're welcome.